OpenStudy (anonymous):

The leg of a right triangle is 3 units and the hypotenuse is 11 units. What is the length, in units, of the other leg of the triangle? 112 units square root 112 units 130 units square root 130 units

OpenStudy (anonymous):

@Music_me

OpenStudy (anonymous):

say somthin ill medal you ;)

OpenStudy (anonymous):

Okie XD

OpenStudy (anonymous):

you know the answer?

OpenStudy (anonymous):

@DJBreezy

OpenStudy (anonymous):

I'm working on it

OpenStudy (anonymous):

pythagorean theorem a^2 + b^2 = c^2 so a = root (c^2 - b^2) a = root ( (11)^2 - (3)^2 ) = ?

OpenStudy (anonymous):

No clue you lost me

OpenStudy (anonymous):

@Music_me

OpenStudy (anonymous):

a^2+b^2=c^2, a=3, c=11. so sqrt of 112=10.58

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

@Music_me

OpenStudy (anonymous):

Yes it is

OpenStudy (anonymous):

WAIT! It's A

OpenStudy (anonymous):

thanks can you help me with 1 more question crazy parson?

OpenStudy (anonymous):

how is it A?

OpenStudy (anonymous):

@mathlover2014

OpenStudy (anonymous):

Sorry, it is B. I read wrong

OpenStudy (anonymous):

ok thanks girly.

OpenStudy (anonymous):

The legs of a right triangle are 6 units and 7 units. What is the length of the hypotenuse? square root 13 units 13 units square root 85 units 85 units

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

YW!! :) HEY!! I AIN'T "GIRLY" :p

OpenStudy (anonymous):

Okie!

OpenStudy (anonymous):

ok your crazy girly

OpenStudy (anonymous):

Yes, tha's better!!

OpenStudy (anonymous):

btw everytime you open a question or help somone, just tag me, i will give you a medal. Just for helping ;)

OpenStudy (anonymous):

Oh, okie :)

OpenStudy (anonymous):

i only have one medal lmao ;)

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

no you didn't have to do that, i was just saying lol, but THANKS ;)

OpenStudy (anonymous):

I always give medals to ppl, just that I sometimes forget to :p

OpenStudy (anonymous):

aww thats nice

OpenStudy (anonymous):

You have the answer yet?

OpenStudy (anonymous):

Thnx :)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Working on it

OpenStudy (anonymous):

np thanks

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @ticklemepickle btw everytime you open a question or help somone, just tag me, i will give you a medal. Just for helping ;) \(\color{blue}{\text{End of Quote}}\) o^_^o

OpenStudy (anonymous):

i tell everyone tht, that helps me

OpenStudy (anonymous):

im nice

OpenStudy (anonymous):

most of times

OpenStudy (anonymous):

lol

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @ticklemepickle The legs of a right triangle are 6 units and 7 units. What is the length of the hypotenuse? square root 13 units 13 units square root 85 units 85 units \(\color{blue}{\text{End of Quote}}\)  apply Pythagorean theorem \[\huge\rm a^2 + b^2 = c^2\] c is the longest side of right triangle which is hypotenuse plug in a nd b value solve for c :-)

OpenStudy (anonymous):

do 6^2 +7 ^2 which equals 85 and then find its square root

OpenStudy (anonymous):

so c?

OpenStudy (anonymous):

why are you doing that @Nnesha

Nnesha (nnesha):

doing what ?

OpenStudy (anonymous):

Yea, it's C

Nnesha (nnesha):

the blue latex ? o^_^o

OpenStudy (anonymous):

@Nnesha @Nnesha @Nnesha @Nnesha @Nnesha @Nnesha @Nnesha @Nnesha

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

eheh it was fun! :D

OpenStudy (anonymous):

YW @ticklemepickle :D

OpenStudy (anonymous):

Do u have any other questions, or is that it?